A bank teller has 54 $5 and $20 bills in her cash drawer. The value of the bills is $780. How many $5 bills are there?
Answer by TimothyLamb(4379)

x = number of $5 bills
y = number of $20 bills
---
x + y = 54
5x + 20y = 780
---
put the system of linear equations into standard form
---
x + y = 54
5x + 20y = 780
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = number of $5 bills = 20
y = number of $20 bills = 34
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php